OpenStudy (anonymous):

The sum of four consecutive integers is 226. What is the THIRD integer?

OpenStudy (anonymous):

here; x+2x+3x+4x=226

OpenStudy (anonymous):

x is any integer

OpenStudy (anonymous):

10x=226 gives x=22.6

OpenStudy (anonymous):

x + x + 1 + x + 2 + x +3 = 4x + 6 = 226

OpenStudy (anonymous):

Subtract 6 from each side of the equation and you end up with 4x = 220. The divide each side by 4

OpenStudy (anonymous):

can anyone help me with MY question THAT I POSTED

OpenStudy (anonymous):

and you get 55. So the four consecutive numbers would be 55, 56, 57, and 58. What is the third number listed?

OpenStudy (anonymous):

57

OpenStudy (anonymous):

lol, is this not the question you posted?

OpenStudy (anonymous):

and yeah that's right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome